OpenStudy (anonymous):

Given the function, f(x) = 1/4x –1, find the value of f(4)

OpenStudy (anonymous):

everwhere you see and x plug in 4 and evluate

OpenStudy (anonymous):

f(4) just means plugging in 4 for the variable, which is x, so the answer is found by following these steps f(x) = 1/4x-1 f(4) = (1/4 * 4) - 1 f(4) = 1-1 = 0 if that x was supposed to be in the denominator then here's the steps f(4) = 1/16 - 1 = 1/16 - 16/16 = - 15/16 but I'm assuming you meant 1/4 * X since the answer comes out cleaner :)

OpenStudy (anonymous):

oh! well what would it be out of these four options, 0, 1, 3/4, -1

OpenStudy (anonymous):

0